write an equation in standard form of the line that passes through the given points- (2,6)(3,8)
Answer by mananth(16946)

x1 y1 x2 y2
2 6 3 8

slope m = (y2-y1)/(x2-x1)
( 8 - 6 )/( 3 - 2 )
( 2 / 1 )
m= 2
The required line passes through both the points. So we select one point and plug it with the slope
Plug value of the slope and point(3,8 ) in
Y = m x + b
8.00 = 2 * 3 + b
8.00 = 6 + b
b= 8 + -6
b= 2
So the equation will be
Y = 2 x + 2
m.ananth@hotmail.ca